OpenStudy (anonymous):

An airplane flew 3 hours with a 40 mph head wind. The return trip with a tail wind of the same speed took 2 hours. Find the speed of the plane in still air.

OpenStudy (ganpat):

head wind is against.. so addition, while tail wind is along so subtraction.. speed = distance/ time 40 = d/ 3 d= 120 miles.. while Return , speed = d/ time s= 120/2 s= 60mph averages speed = 2 (xy)/ (x+y) = 2 (40 * 60)/ 100 = 48mph... this might be answer.. cross check with ur options...

OpenStudy (anonymous):

thanks a lot!!!

OpenStudy (ganpat):

what do option say ??

OpenStudy (anonymous):

oh yea, the options are 180 miles, 200 miles, 220 likes, and 240 miles

OpenStudy (ganpat):

40 = s+ w 60 = s-w so 2s = 100 s= 50 mph.. speed pf plane in still air..